# Poncelet porism in singular cases

### v1

@@ 2.1 Conics in the real plane
Since the parameter of each vertex of the polygonal line depend continuously on the parameter of the first vertex, we get that the only possible limit in each of the cases is point $T$. Thus, a polygonal line inscribed in $\Gamma$ and circumscribed about $\mathcal{C}$ cannot be closed. $\Box$
@@ 2.2 Double intersection point
Let $\Gamma$ and $\mathcal{C}$ be conics in the real plane, such that they have a joint point of order $2$. Then one can choose a coordinate system such that the conics have the following equations:
@@
Proof. We choose the coordinate system as it was done in [FlattoBOOK]: such that the point of joint tangency is placed at the origin and $\mathcal{C}$ is the parabola $y=x^{2}$. $\Box$
@@
which has an ordinary double point at $(\lambda,\mu)=(1/\alpha,0)$, thus its normalization is the following rational curve:
@@
The condition for existence of an $n$-polygon inscribed in $\Gamma$ and circumscribed about $\mathcal{C}$ is that there is a function on the normalized curve (2.3) with zero of order $2n$ at $(\lambda_{1},\mu_{1})=(0,1)$, pole of order $2n$ at $(\lambda_{1},\mu_{1})=(\infty,\infty)$, no other zeros or poles, and having the same values at the two points:
@@ 3 Singular conics
In order to be able to discuss the case when the inscribed conic is not regular, let us observe that the tangent lines to any regular conic are points on the dual of that conic in the dual plane. A degenerate conic in the dual plane consists of the lines from two pencils of lines in the original plane, i.e. of all the lines containing on of the two given points in the original plane. Now, such a degenerate conic allows construction of circumscribed polygons: the sides of the polygon will alternately contain the points of a given pair in the plane.
@@ Acknowledgment
The authors are grateful to Valery Vasilievich Kozlov for the opportunity to learn a lot about various fields of mathematics and mechanics from his books, papers, and lectures, and wish him many happy returns. This research was supported by the Australian Research Council, Discovery Project 190101838 Billiards within quadrics and beyond, the Serbian Ministry of Science, Technological Development and Innovation and the Science Fund of Serbia grant IntegraRS, and the Simons Foundation grant no. 854861.

### v2

@@ 2.1 Conics in the real plane
Since the parameter of each vertex of the polygonal line depends continuously on the parameter of the first vertex, we get that the only possible limit in each of the cases is point $T$. Thus, a polygonal line inscribed in $\Gamma$ and circumscribed about $\mathcal{C}$ cannot be closed. $\Box$
@@ 2.2 Double intersection point
Let $\Gamma$ and $\mathcal{C}$ be conics in the real projective plane, such that they have a joint point of order $2$. Then one can choose an affine chart and a coordinate system such that the conics have the following equations:
@@
Proof. We choose the coordinate system in the projective plane as it was done in [FlattoBOOK]: such that the point of joint tangency is placed at the origin of the affine chart $(x,y)$ and $\mathcal{C}$ is the parabola $y=x^{2}$. With such a choice, the conic $\Gamma$ will have the equation as stated in (2.1). We note that $\alpha\neq 1$, otherwise the order of tangency would be higher than $2$Box
@@
which has an ordinary double point at $(\lambda,\mu)=(-1/\alpha,0)$, thus its normalization is the following rational curve:
@@
The condition for existence of an $n$-polygon inscribed in $\Gamma$ and circumscribed about $\mathcal{C}$ is that there is a function on the normalized curve (2.3) with zero of order $n$ at $(\lambda_{1},\mu_{1})=(0,1)$, pole of order $n$ at $(\lambda_{1},\mu_{1})=(\infty,\infty)$, no other zeros or poles, and having the same values at the two points:
@@ 3 Singular conics
In order to be able to discuss the case when the inscribed conic is not regular, let us observe that the tangent lines to any regular conic are points on the dual of that conic in the dual plane. A degenerate conic in the dual plane consists of the lines from two pencils of lines in the original plane, i.e. of all the lines containing one of the two given points in the original plane. Now, such a degenerate conic allows construction of circumscribed polygons: the sides of the polygon will alternately contain the points of a given pair in the plane.
@@ 3.1 Inscribed conic is singular
    $\mathrm{arctan}\left(\frac{1}{\alpha}\right)\in\left\{\frac{k\pi}{n}\mid 1\leq k<n,(k,n)=1\right\}.$
@@ Acknowledgment
The authors are grateful to Valery Vasilievich Kozlov for the opportunity to learn a lot about various fields of mathematics and mechanics from his books, papers, and lectures, and wish him many happy returns. The authors thank the referees for their helpful comments and suggestions. This research was supported by the Australian Research Council, Discovery Project 190101838 Billiards within quadrics and beyond, the Serbian Ministry of Science, Technological Development and Innovation and the Science Fund of Serbia grant IntegraRS, and the Simons Foundation grant no. 854861.